# A compatible finite element discretisation for moist shallow water equations

### v1

@@ 4.1 Compatible finite element discretisation
We write the weak form of the depth equation by multiplying by a test function $\phi$ from the $\mathbb{V}^{2}$ space and integrating, using integration by parts on the second term. Since $D$ is being transported, we define $D$ on cell edges as the upwind value with respect to the velocity, denoted by $\tilde{D}$. This gives the weak form:
$\int_{\Omega}\phi\frac{\partial D}{\partial t}\text{d}x-\int_{\Omega}D{\bm{u}}\cdot\nabla\phi\text{d}x+\int_{\Gamma}\tilde{D}\llbracket\phi\rrbracket{\bm{u}}\cdot\bm{n}\text{d}S=-\int_{\Omega}\phi\beta_{1}P\text{d}x,\quad\forall\phi\in\mathbb{V}^{2},$ (55)
@@
$\int_{\Omega}\lambda\frac{\partial b}{\partial t}\text{d}x-\int_{\Omega}b\nabla\cdot(\lambda{\bm{u}})\text{d}x+\int_{\Gamma}\tilde{b}\llbracket\lambda{\bm{u}}\rrbracket\text{d}S=\int_{\Omega}\lambda\beta_{2}P\text{d}x,\quad\forall\lambda\in\mathbb{V}^{2},$ (56)
@@ 4.2 Time stepper
The prognostic variables are evolved in time using a semi-implicit quasi-Newton time-stepping scheme. This scheme follows the time stepper used by the Met Office’s models, as described in [27] and [24], and is also the time stepper used in [56] and [52]. The scheme separates the equation into terms that are treated explicitly and terms that are treated semi-implicitly. Each term in the equation is categorised as a forcing term $\mathcal{F}(\chi)$, a transport term $\mathcal{A}(\chi)$, or a physics term $\mathcal{P}(\chi)$, where $\chi=(\bm{u},D,b,q^{k})$ is a single vector encapsulating all prognostic variables. Forcing terms are treated semi-implicitly and transport and physics terms are treated explicitly, with physics happening separately to the dynamics at the end of the time-step. As in, for example, [52] and [56], we use a three-step Runge-Kutta time-stepping procedure (SSPRK3) to timestep the transport terms. Within the Runge-Kutta step, we apply the limiter of Kuzmin [57] as is also described by Cotter and Kuzmin [58] to prevent the generation of new maxima and minima in the transport step.
@@
The next step is an outer iterative loop, which begins with computation of the advecting velocity $\bar{\bm{u}}$ as the average of the velocity at the current time level $\bm{u}^{n}$ and the best guess of the velocity at the next time level $\bm{u}^{n+1}$ : $\bar{\bm{u}}=\frac{1}{2}(\bm{u}^{n}+\bm{u}^{n+1})$. This advecting velocity is then used to explicitly advect the prognostic fields, $\mathcal{A}_{\bar{\bm{u}}}(\chi_{f,e})$, returning $\chi_{T}$:
$\begin{split}\chi_{T}=-\int_{\Omega}\bm{u}_{f,e}\cdot\left(\nabla\times(\bar{\bm{u}}\times\bm{\psi})\right)\text{d}x+\int_{\Gamma}\left(\bm{n}\times\llbracket\bar{\bm{u}}\times\bm{\psi}\rrbracket\right)\cdot\bm{u}_{f,e}^{*}\text{d}S+\int_{\Omega}\nabla\cdot\bm{\psi}\frac{1}{2}(\bm{u}_{f,e}\cdot\bar{\bm{u}})\text{d}x+\int_{\Omega}D_{f,e}{\bar{\bm{u}}}\cdot\nabla\phi\text{d}x\\ -\int_{\Gamma}D_{f,e}^{*}\llbracket\phi\rrbracket{\bar{\bm{u}}}\cdot\bm{n}\text{d}S+\int_{\Omega}b_{f,e}\nabla\cdot(\lambda{\bar{\bm{u}}})\text{d}x-\int_{\Gamma}b_{f,e}^{*}\llbracket\lambda{\bar{\bm{u}}}\rrbracket\text{d}S+\int_{\Omega}q_{f,e}\nabla\cdot(\tau{\bar{\bm{u}}})\text{d}x-\int_{\Gamma}q_{f,e}^{*}\llbracket\tau{\bar{\bm{u}}}\rrbracket\text{d}S\end{split}$ (60)
where $a^{*}$ here symbolises the upwind value of a variable $a$. This step is written in the pseudocode as:
@@
where the $\mathcal{F}$ operator corresponds to the Coriolis and pressure gradient forces. We define a residual $\chi_{r}$ as the difference between this newly-forced state and the best guess for the state at the the next timestep $\chi^{n+1}$ by:
@@
where $\mathcal{J}$ is the Jacobian of $\chi_{r}$ with respect to $\chi$ and the subscript $k$ indicates the iteration index of the Newton loop. To avoid computing the Jacobian at every time step, we instead use a quasi-Newton method and approximate the Jacobian as a simple linear system, which no longer depends on $\chi_{k}^{n+1}$:
@@
$\mathcal{S}\chi^{\prime}=-\chi_{r},$ (67)
where $\chi^{\prime}$ is the increment $\chi_{k+1}^{n+1}-\chi_{k}^{n+1}$.
@@
We begin by linearising the thermal shallow water equations (which do not include physics terms),
@@
$\frac{\partial b^{\prime}}{\partial t}+(\bm{u}^{\prime}\cdot\nabla)b_{ref}=0$ (73)
@@
where the primed quantities represent perturbations from the background state. Then the system given by Equation (67), with $\chi_{r}=(\bm{u}_{r},D_{r},b_{r})$, is:
$\frac{\partial{\bm{u}}^{\prime}}{\partial{t}}+f\bm{\hat{k}}\times\bm{u}^{\prime}+b_{ref}\nabla D^{\prime}+\frac{1}{2}\left[H\nabla b_{ref}+H\nabla b^{\prime}+D^{\prime}\nabla b_{ref}\right]=-\bm{u}_{r},$ (74)
$\frac{\partial D^{\prime}}{\partial t}+\nabla\cdot(\bm{u}^{\prime}H)=-D_{r},$ (75)
$\frac{\partial b^{\prime}}{\partial t}+(\bm{u}^{\prime}\cdot\nabla)b_{ref}=-b_{r}.$ (76)
To solve this we rearrange the buoyancy equation to give an expression for $b^{\prime}$:
$b^{\prime}=-\Delta t(\bm{u}^{\prime}\cdot\nabla)b_{ref}+b_{r},$ (77)
(where we have absorbed a factor of $-\Delta t$ into $b_{r}$) and use this expression for $b^{\prime}$ to solve the system:
$\frac{\partial{\bm{u}}^{\prime}}{\partial{t}}+f\bm{\hat{k}}\times\bm{u}^{\prime}+b_{ref}\nabla D^{\prime}+\frac{1}{2}\left[H\nabla b_{ref}+H\nabla b^{\prime}+D^{\prime}\nabla b_{ref}\right]=-\bm{u}_{r},$ (78)
$\frac{\partial D^{\prime}}{\partial t}+\nabla\cdot(\bm{u}^{\prime}H)=-D_{r},$ (79)
for the updates $(\bm{u}^{\prime},D^{\prime})$ which are then back-substituted to give $b^{\prime}$. This is the same approach as was used for the moist compressible Euler equations in [52].
The solve step then seeks solutions $\bm{u}^{\prime}$ and $D^{\prime}$ by solving a mixed finite element problem, arrived at by multiplying each equation by a test function and integrating:
@@
$\begin{split}\int_{\Omega}{\bm{\psi}}\cdot\frac{\partial{\bm{u}}^{\prime}}{\partial{t}}\text{d}x+\int_{\Omega}{\bm{\psi}}\cdot(f\bm{\hat{k}}\times\bm{u}^{\prime})\text{d}x-\int_{\Omega}\nabla\cdot(\bar{b}{\bm{\psi}})D^{\prime}\text{d}x+\int_{\Gamma}\llbracket\bm{\psi}\bar{b}\rrbracket\langle D^{\prime}\rangle\text{d}S-\frac{1}{2}\int_{\Omega}H\bar{b}\nabla\cdot{\bm{\psi}}\text{d}x-\frac{1}{2}\int_{\Omega}Hb^{\prime}\nabla\cdot{\bm{\psi}}\text{d}x\\ -\frac{1}{2}\int_{\Omega}\bar{b}\nabla\cdot(D^{\prime}{\bm{\psi}})\text{d}x+\frac{1}{2}\int_{\Gamma}\llbracket D^{\prime}{\bm{\psi}}\rrbracket\langle\bar{b}\rangle\text{d}S=-\int_{\Omega}\bm{\psi}\cdot\bm{u}_{r}\text{d}x,\quad\forall\bm{\psi}\in\mathbb{V}^{1},\\ \int_{\Omega}\phi\frac{\partial D^{\prime}}{\partial t}\text{d}x+\int_{\Omega}\phi H\nabla\cdot\bm{u}^{\prime}\text{d}x=\int_{\Omega}\phi D_{r}\text{d}x,\quad\forall\phi\in\mathbb{V}^{2}.\end{split}$ (80)
@@
This is solved using a hybridisation method, which is described in detail in [59, 55] as well as in [52]. The technique involves relaxing the continuity requirements on the normal components of functions in $\mathbb{V}^{1}$ and enforcing continuity through Lagrange multipliers as part of the solution formulation. We then back-substitute using these $(\bm{u}^{\prime},D^{\prime})$ values to get a value for $b^{\prime}$. In the moist convective formulation there is no prognostic buoyancy field, and so we linearise the usual shallow water equations rather than the thermal shallow water equations. Again we arrive at a system for $\bm{u}^{\prime}$ and $D^{\prime}$ which we solve by writing a mixed finite element problem and applying a hybridisation method. These increments $\chi^{\prime}=(\bm{u}^{\prime},D^{\prime},b^{\prime})$ (or $\chi^{\prime}=(\bm{u}^{\prime},D^{\prime})$ in the moist convective model) are added to the current guess to update $\chi^{n+1}$ and the iteration begins again.
@@
9:         Solve: $\mathcal{S}(\chi^{\prime})=\chi_{r}$ for $\chi^{\prime}$
10:         Increment: $\chi^{n+1}=\chi^{n+1}+\chi^{\prime}$
@@ 5 Results
In this section we give details of three test cases in each moist shallow water formulation, along with results. Our model is written using Gusto, the dynamical core toolkit built on the Firedrake finite element library. Firedrake provides automated code generation for the solution of partial differential equations using the finite element method [60].
For the Earth-related physical constants for all tests we take rotation rate of the Earth $\Omega=7.292\times 10^{-5}\text{s}^{-1}$, acceleration due to gravity $g=9.80616\text{ m }\text{s}^{-2}$ and radius of the Earth $R=6371220\text{ m}$. As described in Zerroukat and Allen [7], in all our tests the initial vapour is dependent on the saturation, and is chosen to be near saturation so that the physics is activated easily and cloud and rain are produced without having to run the model for a long time. The scaling factor $q_{0}$ that appears in the saturation function $q_{sat}$ is chosen so that the maximum value of the initial vapour is around the atmospheric value of 0.02. For both the steady state test and the flow over a mountain test that means a value for $q_{0}$ of 0.007, and a value of 0.0027 for the unstable jet test. Rain, when it is produced, is not advected with the flow but instead accumulates where it forms.
@@
with $u_{0}=20\text{ m s}^{-1}$, $\omega=\left(\Omega au_{0}+\frac{u_{0}^{2}}{2}\right)$, $\sigma=\omega/10$, $\Phi_{0}=3\times 10^{4}\text{ m s}^{-2}$, and $\theta_{0}=\epsilon\Phi_{0}^{2}$ with $\epsilon=1/300$.
@@ 5.1 Moist Steady State Geostrophic Flow
The first test is a modification of test 2 of Williamson et al. [25], and is based on the steady state test of Zerroukat and Allen [7] (written there in terms of geopotential rather than depth, and potential temperature rather than buoyancy) [^3] Note the sign change in Equation (87) relative to the potential temperature equation in Zerroukat and Allen [7] to ensure balanced initial conditions.. The test consists of zonally balanced flow on a sphere with no topography. The initial conditions are solutions to the thermal shallow water equations and so are a slight modification of the original Williamson et al. initial conditions to account for the extra terms in the thermal system. Because the test is initialised with a steady state, the expectation is that the solution should remain at the initial state throughout the simulation, and thus the metric for evaluating the results is how well the final state compares to the initial state. The model is initialised with water vapour $q_{v}$ either at or close to saturation everywhere and no cloud or rain. The initial conditions for both of the models with buoyancy (moist thermal and moist convective thermal) are (with $(\lambda,\phi)$ representing spherical (longitude, latitude) coordinates):
@@
where $u_{0}=20\text{ m s}^{-1}$, $\omega=\left(\Omega au_{0}+\frac{u_{0}^{2}}{2}\right)$, $\sigma=\omega/10$, $\Phi_{0}=3\times 10^{4}\text{ m s}^{-2}$, and $\theta_{0}=\epsilon\Phi_{0}^{2}$ with $\epsilon=1/300$. The background depth $H$ is $\Phi_{0}/g$. $\xi$ is a parameter that controls how near to saturation the initial vapour is. When $\xi=0$ the vapour is initialised with the initial saturation values, and when $\xi=10^{-3}$ vapour is initialised just below saturation.

### v2

@@ 4.1 Compatible finite element discretisation
We write the weak form of the depth equation (Equation (16)) by multiplying by a test function $\phi$ from the $\mathbb{V}^{2}$ space and integrating, using integration by parts on the second term. Since $D$ is being transported, we define $D$ on cell edges as the upwind value with respect to the velocity, denoted by $\tilde{D}$. This gives the weak form:
$\int_{\Omega}\phi\frac{\partial D}{\partial t}\text{d}x-\int_{\Omega}D{\bm{u}}\cdot\nabla\phi\text{d}x+\int_{\Gamma}\tilde{D}\llbracket\phi\rrbracket{\bm{u}}\cdot\bm{n}\text{d}S=\int_{\Omega}\phi\beta_{1}S_{D}\text{d}x,\quad\forall\phi\in\mathbb{V}^{2},$ (47)
@@
$\int_{\Omega}\lambda\frac{\partial b}{\partial t}\text{d}x-\int_{\Omega}b\nabla\cdot(\lambda{\bm{u}})\text{d}x+\int_{\Gamma}\tilde{b}\llbracket\lambda{\bm{u}}\rrbracket\text{d}S=\int_{\Omega}\lambda\beta_{2}S_{b}\text{d}x,\quad\forall\lambda\in\mathbb{V}^{2},$ (48)
@@ 4.2 Time stepper
The prognostic variables are evolved in time using a semi-implicit quasi-Newton time-stepping scheme. This scheme follows the time stepper used by the Met Office’s models, as described in [27] and [24], and is also the time stepper used in [56] and [52]. The scheme separates the equation into terms that are treated explicitly and terms that are treated semi-implicitly. Each term in the equation is categorised as a forcing term $\mathcal{F}(\chi)$, a transport term $\mathcal{A}(\chi)$, or a physics term $\mathcal{P}(\chi)$, where $\chi=(\bm{u},D,b,q^{k})$ is a single vector encapsulating all prognostic variables. Forcing terms are treated semi-implicitly and transport and physics terms are treated explicitly, with physics happening separately to the dynamics at the end of the time-step.
@@
The next step is an outer iterative loop, which begins with computation of the advecting velocity $\bar{\bm{u}}$ as the average of the velocity at the current time level $\bm{u}^{n}$ and the best guess of the velocity at the next time level $\bm{u}^{n+1}$ : $\bar{\bm{u}}=\frac{1}{2}(\bm{u}^{n}+\bm{u}^{n+1})$. This advecting velocity is then used to explicitly advect the prognostic fields, $\mathcal{A}_{\bar{\bm{u}}}(\chi_{f,e})$, returning $\chi_{T}$. For the transport of $\bm{u}$, the equation is written in vector-invariant form and timestepped using the implicit midpoint method. All other fields are transported using an upwind discontinuous Galerkin scheme with a three-step Runge-Kutta timestepping procedure (SSPRK3), as outlined in, for example [52] and [56]. Within the Runge-Kutta step, we apply the limiter of Kuzmin [57] as is also described by Cotter and Kuzmin [58] to prevent the generation of new maxima and minima in the transport step.
This step is written in the pseudocode as:
@@
where the $\mathcal{F}$ operator corresponds to the Coriolis and pressure gradient forces. We define a residual $\chi_{r}$ as the difference between this newly-forced state and the best guess for the state at the next timestep $\chi^{n+1}$ by:
@@
where $\mathcal{J}$ is the Jacobian of $\chi_{r}$ with respect to $\chi^{n+1}$ and the subscript $k$ indicates the iteration index of the Newton loop. To avoid computing the Jacobian at every time step, we instead use a quasi-Newton method and approximate the Jacobian as a simple linear system, which no longer depends on $\chi_{k}^{n+1}$:
@@
$\mathcal{S}\delta\chi=-\chi_{r},$ (58)
where $\delta\chi$ is the increment $\chi_{k+1}^{n+1}-\chi_{k}^{n+1}$.
@@
To derive our linear operator $\mathcal{S}$, we begin by linearising the thermal shallow water equations (which do not include physics terms),
@@
$\frac{\partial b^{\prime}}{\partial t}+(\bm{u}^{\prime}\cdot\nabla)b_{ref}=0,$ (64)
@@
where the primed quantities represent perturbations from the background state.
Using the operator for this linear system to approximate the Jacobian, we obtain the matrix-vector problem corresponding to Equation (58):
@@
$\begin{split}\int_{\Omega}{\bm{\psi}}\cdot\delta\bm{u}\text{d}x+\Delta t\int_{\Omega}{\bm{\psi}}\cdot(f\bm{\hat{k}}\times\delta\bm{u})\text{d}x-\Delta t\int_{\Omega}\nabla\cdot(\bar{b}{\bm{\psi}})\delta D\text{d}x+\Delta t\int_{\Gamma}\llbracket\bm{\psi}\bar{b}\rrbracket\langle\delta D\rangle\text{d}S-\frac{\Delta t}{2}\int_{\Omega}H\bar{b}\nabla\cdot{\bm{\psi}}\text{d}x\\ -\frac{\Delta t}{2}\int_{\Omega}H\delta b\nabla\cdot{\bm{\psi}}\text{d}x-\frac{\Delta t}{2}\int_{\Omega}\bar{b}\nabla\cdot(\delta D{\bm{\psi}})\text{d}x+\frac{\Delta t}{2}\int_{\Gamma}\llbracket\delta D{\bm{\psi}}\rrbracket\langle\bar{b}\rangle\text{d}S=-\int_{\Omega}\bm{\psi}\cdot\bm{u}_{r}\text{d}x,\quad\forall\bm{\psi}\in\mathbb{V}^{1},\\ \int_{\Omega}\phi\delta D\text{d}x+\Delta t\int_{\Omega}\phi H\nabla\cdot\delta\bm{u}\text{d}x=-\int_{\Omega}\phi D_{r}\text{d}x,\quad\forall\phi\in\mathbb{V}^{2},\\ \int_{\Omega}\lambda\delta b\text{d}x-\Delta t\int_{\Omega}b_{ref}\nabla\cdot(\lambda{\delta\bm{u}})\text{d}x+\Delta t\int_{\Gamma}\tilde{b}_{ref}\llbracket\lambda{\delta\bm{u}}\rrbracket\text{d}S=-\int_{\Omega}\lambda b_{r}\text{d}x,\quad\forall\lambda\in\mathbb{V}^{2},\end{split}$ (65)
where the increment we are solving for, $\delta\chi$, is given by $\delta\chi=(\delta\bm{u},\delta D,\delta b)$. We solve this system by eliminating the buoyancy equation via:
$\delta b=-\Delta t(\delta\bm{u}\cdot\nabla)b_{ref}+b_{r},$ (66)
@@
to give a reduced system for $\bm{u}$ and $D$. This reduced system is solved using a hybridisation method, which is described in detail in [59, 55] as well as in [52]. The technique involves relaxing the continuity requirements on the normal components of functions in $\mathbb{V}^{1}$ and enforcing continuity through Lagrange multipliers as part of the solution formulation. We then back-substitute using these $(\delta\bm{u},\delta D)$ values to get a value for $\delta b$. In the moist convective formulation there is no prognostic buoyancy field, so our operator $\mathcal{S}$ is derived by linearising the standard shallow water equations, resulting in a system for $\delta\bm{u}$ and $\delta D$ which is solved in the same way as the reduced system above. These increments $\delta\chi=(\delta\bm{u},\delta D,\delta b)$ (or $\delta\chi=(\delta\bm{u},\delta D)$ in the moist convective model) are added to the current guess to update $\chi^{n+1}$ and the iteration begins again.
@@
9:         Solve: $\mathcal{S}(\delta\chi)=\chi_{r}$ for $\delta\chi$
10:         Increment: $\chi^{n+1}=\chi^{n+1}+\delta\chi$
@@ 5 Results
In this section we give details of three test cases in each moist shallow water formulation, along with results. Our model is written using Gusto, the dynamical core toolkit built on the Firedrake finite element library. Firedrake provides automated code generation for the solution of partial differential equations using the finite element method [60]. Code is automatically generated using the Unified Form Language [61] and equations are provided to PETSc [62], which provides direct access to runtime configurable iterative solvers and preconditioners. The hybridisation of the linear system in the linear solve stage is implemented using the Slate framework [55], which performs local elimination and recovery options. The reduced equation for the trace variables is numerically inverted using the conjugate gradient method and PETSc’s GAMG preconditioner.
Our first step to verify our model was to run Test 2 from [7] using the moist thermal formulation, with the initial conditions as given there. We found that it was necessary to reproduce as closely as we could the resolution and timestep in [7] in order to see the same amount of cloud and rain. We also found that the amount of cloud and rain were both sensitive too to the use of the limiters in the transport step, and that more of both species were generated when no limiters were used. This experiment, while establishing confidence in our model, also demonstrated the sensitivity of the cloud field to aspects of the model and various modelling choices. (These results are not presented here.)
For the Earth-related physical constants for all tests we take rotation rate of the Earth $\Omega=7.292\times 10^{-5}\text{s}^{-1}$ where $f=2\Omega\cos{\phi}$ with $\phi$ the latitude, acceleration due to gravity $g=9.80616\text{ m }\text{s}^{-2}$ and radius of the Earth $R=6371220\text{ m}$. As described in Zerroukat and Allen [7], in all our tests the initial vapour is dependent on the saturation, and is chosen to be near saturation so that the physics is activated easily and cloud and rain are produced without having to run the model for a long time. The scaling factor $q_{0}$ that appears in the saturation function $q_{sat}$ is chosen so that the maximum value of the initial vapour is around the atmospheric value of 0.02. For both the steady state test and the flow over a mountain test that means a value for $q_{0}$ of 0.007, and a value of 0.0027 for the unstable jet test. Rain, when it is produced, is not advected with the flow but instead accumulates where it forms.
@@
with $u_{0}=20\text{ m s}^{-1}$, $\omega=\left(\Omega au_{0}+\frac{u_{0}^{2}}{2}\right)$, $\sigma=\omega/10$, $\Phi_{0}=3\times 10^{4}\text{ m}^{2}\text{ s}^{-2}$, and $\theta_{0}=\epsilon\Phi_{0}^{2}$ with $\epsilon=1/300$.
@@ 5.1 Moist Steady State Geostrophic Flow
The first test is a modification of test 2 of Williamson et al. [25], and is based on the steady state test of Zerroukat and Allen [7] (written there in terms of geopotential rather than depth, and potential temperature rather than buoyancy) [^3] Note the sign change in Equation (73) relative to the potential temperature equation in Zerroukat and Allen [7] to ensure balanced initial conditions.. The test consists of zonally balanced flow on a sphere with no topography. The initial conditions are solutions to the thermal shallow water equations and so are a slight modification of the original Williamson et al. [25] initial conditions to account for the extra terms in the thermal system. Because the test is initialised with a steady state, the expectation is that the solution should remain at the initial state throughout the simulation, and thus the metric for evaluating the results is how well the final state compares to the initial state. The model is initialised with water vapour $q_{v}$ either at or close to saturation everywhere and no cloud or rain. The initial conditions for both of the models with buoyancy (moist thermal and moist convective thermal) are (with $(\lambda,\phi)$ representing spherical (longitude, latitude) coordinates):
@@
where $u_{0}=20\text{ m s}^{-1}$, $\omega=\left(\Omega Ru_{0}+\frac{u_{0}^{2}}{2}\right)$, $\sigma=\omega/10$, $\Phi_{0}=3\times 10^{4}\text{ m}^{2}\text{ s}^{-2}$, and $\theta_{0}=\epsilon\Phi_{0}^{2}$ with $\epsilon=1/300$. The background depth $H$ is $\Phi_{0}/g$. $\xi$ is a parameter that controls how near to saturation the initial vapour is. When $\xi=0$ the vapour is initialised with the initial saturation values, and when $\xi=10^{-3}$ vapour is initialised just below saturation.